OpenStudy (anonymous):

|-13| + 42 = 100y - |-32| - 99y solve the equation. need help.

OpenStudy (anonymous):

What do absolute value bars do to the numbers inside of them?

OpenStudy (anonymous):

makes them whole? that part i know it's the rest of the problem i get lost lol

OpenStudy (anonymous):

Yes, it makes them positive. So we can already simplify all the numbers between the absolute value bars. That will yield: \[13 + 42 = 100y - 32 - 99y\] Do you know what to do from here?

OpenStudy (anonymous):

subtract y?

OpenStudy (anonymous):

Actually, you might want to look for like-terms first. 13 and 42 are like terms. Once we add them together we get: \[55 = 100y-32-99y\] Also, we can combine 100 and -99y. That gives us 1 y. Now we have: \[55 = y - 32\]Now we need to move all the terms containing y to the right, all other terms to the left. Can you tell me how to move 32 to the left?

OpenStudy (anonymous):

subtract from 55?

OpenStudy (anonymous):

an thanks you helped me alot lol figuring out what i did wrong

OpenStudy (anonymous):

actually, remember, what you do to one side of the equation you must do to the other. That gives you a hint as to what you need to do to move 32 to the left.

OpenStudy (anonymous):

\[55=y - 32\] Close, but the 55 is fine on the left side of the equation. We can leave it alone. However, we need to do something with the 32 to move it over to the left.. like @juan_9292 said. And, you're welcome! I'm glad I can help.

OpenStudy (anonymous):

okay now i'm confused lol

OpenStudy (anonymous):

do we combine like terms again?

OpenStudy (anonymous):

Okay, we want to get the "y" term alone, right? So that means we need to get rid of the "-32". What can we add to -32 to make it equal 0?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

Yes! And since what you do to one side of the equation you must do to the other, we need to add 32 to the left side of the equation as well. \[55=y-32\]\[55 + 32 = y -32 +32\] \[87 = y + 0\]\[y = 87\] And there's your answer! :D Does that make sense?

OpenStudy (anonymous):

we did something wrong lol my multiple choice say 3 -3 61 -61

OpenStudy (anonymous):

Hahah oh dang! Let me go back and check my work, hold on.

OpenStudy (anonymous):

I'm pretty sure that's the right answer for the problem you provided, though. D: Are you sure you typed in the right problem?

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

no

OpenStudy (anonymous):

crap lmao

OpenStudy (anonymous):

it's not 42it was 4 to the power of 2 lol

OpenStudy (anonymous):

i'm sorry lol

OpenStudy (anonymous):

Oh, hahaha! That's okay, we all make mistakes. Let me try to solve it, one second.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Okay, it's essentially the same problem.\[|-13| + 4^2 = 100y - |-32| - 99y\] We talked about before how to simplify absolute values.\[13 + 4^2 = 100y - 32 - 99y\] Okay, the order of operations are PEMDAS, as I'm sure you know. Since we don't have any parenthesis, we can simplify the exponents. Whats 4^2?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Right! \[13+16=100y-32-99y\] So can you combine like terms?

OpenStudy (anonymous):

13+16?

OpenStudy (anonymous):

Yes, so what is 13 + 16?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

Correct, so now we have: \[29 = 100y-32-99y\] Can we combine any other like terms? I'll give you a hint, you can see them easily if I rearrange the equation like this: \[29 = 100y-99y -32\]

OpenStudy (anonymous):

99 and 32?

OpenStudy (anonymous):

We can't combine -99y and -32 because -99 has a y attached to it, and -32 doesn't. Do you see any two terms that both have a y attached to them that we could easily combine?

OpenStudy (anonymous):

oh lol 100y and 99y

OpenStudy (anonymous):

Yes! So if we compute 100y - 99y, what do we get?

OpenStudy (anonymous):

1y?

OpenStudy (anonymous):

Correct! So now we have the equation:\[29 = y - 32\] How can we get the "y" alone by itself? In other words, how can we move the -32 to the left side of the equation?

OpenStudy (anonymous):

minus 32 with 32

OpenStudy (anonymous):

or wiyth 29?

OpenStudy (anonymous):

oh it will be -3?

OpenStudy (anonymous):

No, you were right the first time, with 32. However, we need to do the opposite operation in order to cancel it out. It's already subtracting from y, so we need to add it. Does that make sense?

OpenStudy (anonymous):

that's since you want y on its own, all alone

OpenStudy (anonymous):

so 29=64?

OpenStudy (anonymous):

or 61=64?

OpenStudy (anonymous):

That's close! Let's take a look at our equation again: \[29 = y -32\] We want y all alone on one side, so we need to add 32 to both sides. When we compute -32 + 32, what do we get?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Yes! \[29=y-32\] Like we figured out before, we need to add 32 to both sides.\[29 (+32) = y -32 (+32)\] You correctly said that -32+32 is 0 so:\[29(+32) = y - 0\] y - 0 is y, so we don't even have to write the 0.\[29(+32) = y \] You're almost there! Whats 29 + 32?

OpenStudy (anonymous):

61

OpenStudy (anonymous):

Yes, so \[61 = y\] We can flip it around so it's:\[y=61\] And there's your answer! Does that make sense? Make sure you understand.

OpenStudy (anonymous):

yes thank you (:

OpenStudy (anonymous):

You're welcome! :D

OpenStudy (anonymous):

@friedch1cken i need your help i did the problem |-17| + 52 = 20y - |-35| - 19y wrong i got y=-42 how i got that answered was i did -17+25=20y-35-19y so i subtracted -19 y from the like terms with the y and got -17+25=y so i subtracted -25 from -17 and 25 and got -42

OpenStudy (anonymous):

but the multiple choice say -7 7 -77 77

OpenStudy (anonymous):

well lets see, you said that the problem was |-17| +52 = 20y - |-35|- 19y hmm I got confused as to how you did your work; but lets do this from scratch ok? Or would you like me to point out a mistake?

OpenStudy (anonymous):

point out mistake lol

OpenStudy (anonymous):

lol, ok, well first mistake I see that you forgot to get the absolute value of the number inside

OpenStudy (anonymous):

the absolute value of -17 or |-17| is actually 17. and so on

OpenStudy (anonymous):

oh that's what i did wrong lol

OpenStudy (anonymous):

yea, and also, I see that you changed 52 into 25 when you were doing your work, perhaps that made you get such a high number.

OpenStudy (anonymous):

i kept it negative lol *slaps self*

OpenStudy (anonymous):

I think she meant 5^2.

OpenStudy (anonymous):

yeah i meant 5^2 lol happened again when i typed

OpenStudy (anonymous):

oh, coincides with her work; yea, nevermind about that then :p

OpenStudy (anonymous):

Good thing we caught it early :D

OpenStudy (anonymous):

thank you both ^_^

OpenStudy (anonymous):

no problemo :)

OpenStudy (anonymous):

Okay, let me know what you get for an answer :D

OpenStudy (anonymous):

lol, I see why you said "Good thing we caught it early"